Two independent events have probabilities 0.4 and 0.5. Probability both occur?

Two independent events have probabilities 0.4 and 0.5. Probability both occur?

Approach: Multiply probabilities.

0.2. 0.4 * 0.5 = 0.2.

Follow-up: Probability at least one occurs?

Key concepts: joint probability.